OpenStudy (anonymous):

how do you simplify 16 plus or minus the square root of 32 all over 14

OpenStudy (anonymous):

\[62=16\times 2\] so \[\sqrt{32}=\sqrt{16}\sqrt2=4\sqrt2\] is a start

OpenStudy (anonymous):

you have \[\frac{16\pm\sqrt{32}}{14}=\frac{16\pm4\sqrt2}{14}=\frac{2(8+2\sqrt2)}{14}\] then cancel the 2 top and bottom

OpenStudy (anonymous):

so the final result is ?